OpenStudy (anonymous):

Evaluate the expression 3.14(a2 + ab) when a = 4 and b = 2. (Input decimals only, such as 12.71, as the answer.)

OpenStudy (taramgrant0543664):

All you have to do is put 4 in for a and 2 in for b, and put it into your calculator

OpenStudy (anonymous):

can you help more?

OpenStudy (anonymous):

thanks c:

OpenStudy (anonymous):

i got 67.12?

OpenStudy (anonymous):

@taramgrant0543664

OpenStudy (taramgrant0543664):

Is it 3.14((2xa)+(ab)) or 3.14((a^2)+(ab))?

OpenStudy (anonymous):

3.14(a2 + ab)

OpenStudy (taramgrant0543664):

So what I got then is 50.24 3.14(2(4)+(4)(2))=3.14(8+8)=3.14(16)=50.24

OpenStudy (anonymous):

im kind of confused

OpenStudy (anonymous):

thank you @taramgrant0543664 for your help